OpenStudy (anonymous):

use logathrimic differentiation to differentiate the following: y = 3 ^ tanx thanks for any help

OpenStudy (amistre64):

log and diff ... right?

OpenStudy (anonymous):

Take the natural log of both sides and use properties of logs to get \[\ln(y)=\tan(x)\ln(3)\] ln(3) is a constant, so when you differentiate, you can treat it like any other constant. \[\ln(y)'=(\ln(3)\tan(x))'\] Pull out the ln(3)\[\frac{1}{y}=\ln(3)(\tan(x))'\]\[\frac{1}{y}=\ln(3)\sec^2(x)\] Flip it and you're done.

OpenStudy (amistre64):

ln (y = 3 ^ tanx) lny = ln 3 ^ tanx lny = tanx ln 3 y'/y = tan'x ln 3 + tanx ln' 3 y'/y = sec^2 x ln 3 y' = y(sec^2 x ln 3)

OpenStudy (anonymous):

ln3 is a constant, so you don't need to use chain rule, I think. It should go to 0 anyway.

OpenStudy (amistre64):

since constants go to zero; the chain rule works on them and they zero out; and your left with the so called "constant" rule

OpenStudy (anonymous):

i didnt know whether ln 3 went to 1/3..like most ln derivatives do..thanks so much for the quick reply

OpenStudy (amistre64):

[2x]' = 2x'+2'x [2x]' = 2+0x [2x]' = 2

OpenStudy (anonymous):

Right, what I meant by the going to 0 part. Where is the y coming from?

OpenStudy (amistre64):

ln3 is just a constant; no need to try to make it a variable ;)

OpenStudy (amistre64):

implicit differentation; y -> y'

OpenStudy (anonymous):

Oh, right, duh. Thanks! Blonde moment.

OpenStudy (amistre64):

ln(y) chains to; 1/y * y'

OpenStudy (anonymous):

thanks so much for the help

OpenStudy (amistre64):

yw